# Challenges in Testing Large Language Model Based Software: A Faceted Taxonomy

### v1

@@ Abstract.
Large Language Models (LLMs) and Multi-Agent LLMs (MALLMs) introduce non-determinism unlike traditional or machine learning software, requiring new approaches to verifying correctness beyond simple output comparisons or statistical accuracy over test datasets.
This paper presents a taxonomy for LLM test case design, informed by both the research literature, our experience, and open-source tools that represent the state of practice. We identify key variation points that impact test correctness and highlight open challenges that the research, industry, and open-source communities must address as LLMs become integral to software systems.
Our taxonomy defines four facets of LLM test case design, addressing ambiguity in both inputs and outputs while establishing best practices. It distinguishes variability in goals, the system under test, and inputs, and introduces two key oracle types: atomic and aggregated. Our mapping indicates that current tools insufficiently account for these variability points, highlighting the need for closer collaboration between academia and practitioners to improve the reliability and reproducibility of LLM testing.
@@ 1. Introduction
To address this gap, Section 2 introduces our faceted taxonomy (Usman et al., 2017) for LLM test cases, categorizing key variation points that impact evaluation correctness, including the distinction between Atomic and Aggregated Oracles. Unlike prior work–(Xiao et al., 2024) on adversarial robustness and (Hudson et al., 2024) on broad LLM testing taxonomies–our framework organizes test case design across multiple facets, extending beyond specific testing foci and high-level categorization. Section 3 then maps current practices onto this taxonomy, assessing coverage across dimensions. Finally, Section 4 identifies key challenges for ensuring correctness in LLM-based software.
@@ 2.4. Inputs
Addressing input variability requires systematic coverage strategies across syntactic and semantic dimensions. This enhances reliability and robustness in LLM-based systems while clarifying failure points linked to input ambiguities.
@@
3. Mapping the Taxonomy
@@
We identified relevant GitHub repositories by searching for “LLM Testing” and “LLM Evaluation”, selecting frameworks that specifically target LLM applications (not foundational models) and provide dedicated testing interfaces. The three selected projects, chosen based on GitHub popularity (stars) and distinct testing features, are Promptfoo (Team, 2024b) (5.6k stars), DeepEval (Team, 2024a) (5.1k stars), and Giskard (Giskard Team, 2024) (4.3k stars).
@@
Facets Promptfoo DeepEval Giskard
SUT SUT variations can be compared using visualization features. SUT variations across versions can be analyzed using visualization features to detect regressions. SUT variations are less emphasized but can be still compared via external platform integration.
Goal Testing goal can be set with one or more properties with varying number of inputs. Testing goal can be set with one or more properties for each test input. Testing goal is defined as one or more properties across a set of inputs.
Oracles Oracles are implemented using deterministic and model-based predefined metrics, allowing weighted metric combinations. Oracles are handled mostly using model-based metrics, including conversational/multi-modal support. Oracles are typically implemented as dataset-wise scores via built-in metric functions across multiple inputs.
Inputs Inputs with semantic variations can be synthesized with an separate LLM agent. Inputs with semantic variations can be synthesized with an separate LLM agent and with various resources (e.g., documents). Inputs can be transformed using various predefined syntactic transformation methods.
@@
All tools accommodate variability across the three SUT sub-facets, though their capabilities differ slightly. Each allows users to specify the Component under test (e.g., a prompt template or agent), select the Model(s) (e.g., GPT-4o, Claude 3.7 Sonnet), and adjust Configurations (e.g., temperature, token limits) that influence runtime behavior. Notably, Promptfoo enables “unfolded” comparisons, allowing testers to assess differences between prompt versions, model choices, or configuration settings. Similarly, DeepEval provides visualization tools but primarily focuses on detecting regressions over time. For cross-version SUT comparisons, Giskard integrates with external commercial platforms like Weights and Biases (Weights & Biases, 2025).
@@
An ideal test case is defined with a clear and specific testing goal, typically composed of one or more relevant properties. In Promptfoo and DeepEval, a test case consists of a single input paired with an expected output, evaluated against one or more property-checking metrics. In contrast, Giskard integrates into an ML evaluation pipeline, running tests on entire datasets and producing dataset-wide scores. Some LLM testing sub-goals–such as assessing robustness against prompt injection attacks–are best evaluated using a set of targeted inputs. While Promptfoo and DeepEval can execute multiple test cases in a single evaluation, they lack explicit support for linking evaluations to a coherent test goal. Conversely, Giskard lacks the flexibility to configure individual test cases separately when needed.
@@
Evaluations typically yield either a continuous score (e.g., similarity to a ground-truth output on a 0-1 scale) or a binary pass/fail result, with continuous scores often thresholded into binary outcomes. Giskard operates at the dataset level, aggregating individual assessments (e.g., reporting a 70% pass rate for politeness). However, LLM outputs can vary across repeated runs, and current tools lack aggregation methods to handle multiple evaluations of the same input within the same system under test (SUT). Instead, each run is treated as an independent outcome, ignoring stochastic variability.

### v2

@@ Abstract.
Large Language Models (LLMs) and Multi-Agent LLMs (MALLMs) introduce non-determinism unlike traditional or machine learning software, requiring new approaches to verifying correctness beyond simple output comparisons or statistical accuracy over test datasets. This paper presents a taxonomy for LLM test case design, informed by research literature and our experience. Each facet is exemplified, and we conduct an LLM-assisted analysis of six open-source testing frameworks, perform a sensitivity study of an agent-based system across different model configurations, and provide working examples contrasting atomic and aggregated test cases. We identify key variation points that impact test correctness and highlight open challenges that the research, industry, and open-source communities must address as LLMs become integral to software systems. Our taxonomy defines four facets of LLM test case design, addressing ambiguity in both inputs and outputs while establishing best practices. It distinguishes variability in goals, the system under test, and inputs, and introduces two key oracle types: atomic and aggregated. Our findings reveal that current tools treat test executions as isolated events, lack explicit aggregation mechanisms, and inadequately capture variability across model versions, configurations, and repeated runs. This highlights the need for viewing correctness as a distribution of outcomes rather than a binary property, requiring closer collaboration between academia and practitioners to establish mature, variability-aware testing methodologies.
@@ 1. Introduction
This article extends our previous work (Dobslaw et al., 2025) by presenting a faceted taxonomy for LLM test case design that organizes testing concerns across four dimensions: Software Under Test, Goal, Oracles, and Inputs. Central to this taxonomy is the distinction between atomic oracles, which evaluate single executions under the assumption of determinism, and aggregated oracles, which assess correctness across multiple runs to account for stochastic behavior. We validate the taxonomy through comprehensive empirical investigations that combine multiple methods: concrete examples demonstrating each facet, LLM-assisted analysis of six open-source testing frameworks (Opik, DeepEval, RAGAs, Promptfoo, Phoenix, and Giskard), a sensitivity study comparing different model configurations (GPT-3.5-turbo vs GPT-4) in the DroidAgent agent-based system, and working implementations contrasting atomic and aggregated test approaches.
Our findings reveal gaps across multiple dimensions of current testing tooling and practice. First, analysis of the six frameworks shows that tools predominantly treat each test execution as an isolated event, relying on atomic oracles that presume deterministic outcomes. Second, while some tools aggregate results at the dataset level (e.g., reporting pass rates), none provide explicit mechanisms for handling multiple evaluations of the same input within a single system configuration–ignoring the stochastic variability inherent to LLMs. Third, the sensitivity study demonstrates that even minor configuration changes (model selection, temperature settings) can dramatically alter system behavior, yet current tools lack systematic support for capturing and comparing such variability across model versions and configurations. These observations collectively challenge the assumption that correctness can be assessed through isolated executions and reveal that mature LLM testing requires viewing correctness as a distribution of outcomes rather than a binary property. This conceptual shift has implications not only for tool design and aggregation strategies but also for how testing responsibilities may need to extend beyond traditional software testers to include domain experts who possess the contextual knowledge necessary for validation in specialized domains.
@@
The structure of the article is as follows. Section 2 offers background and motivates the need for a taxonomy and structured approach to the testing of LLM-based software. Section 3 introduces our faceted taxonomy (Usman et al., 2017) for LLM test cases, categorizing key variation points that impact evaluation correctness, including the distinction between Atomic and Aggregated Oracles to address the identified gap. Unlike prior work–(Xiao et al., 2024) on adversarial robustness and (Hudson et al., 2024) on broad LLM testing taxonomies–our framework organizes test case design across multiple facets, extending beyond specific testing foci and high-level categorization. Section 4 presents our empirical investigations, including tool mapping, LLM-assisted analysis, and sensitivity studies. Finally, Section 5 discusses implications for the field and identifies key challenges for ensuring correctness in LLM-based software toward 2030.
2. Background and Motivation
The proposed taxonomy for LLM test cases is designed to support not only the traditional testing of finalized LLM-based systems but also to play a central role throughout the entire development, maintenance, and evolution phases. Given the inherent non-determinism, flexible behavior, and the vast and often unpredictable output space of LLMs, we believe that test cases will be an essential part of the development process of any LLM-based systems, far more so than for the traditional software systems.
While the emphasis on the need for test cases throughout the development lifecycle strongly reminds us of Test-Drive Development (TDD), we argue that LLM-based systems would take on a more dynamic and iterative process than that of TDD as we know of. In conventional software development, TDD follows a relatively linear cycle: developers write a test case which declaratively specifies what the expected program behavior should be, implement code to pass the test, and maintain a stable, “green” test suite–unless the specification changes. However, in LLM-based systems, the development process is less deterministic. Developers often begin with only a vague understanding of “what to test”, especially in domains with complex requirements (e.g., medical chatbots). The high degree of causal uncertainty–the inability to predict how an LLM will interpret and follow a given prompt–means that developers cannot statically reason about expected behaviors. Unlike traditional code, where the flow of execution can often be understood without running it, at least on lower levels, LLM behaviors must be explored through dynamic analysis, observing actual outputs to identify patterns, strengths, and failures.
This leads to a more fluid form of TDD, where both the Software Under Test (SUT) and the test cases themselves evolve together. Developers iteratively refine prompts, configurations, and even the underlying models, all while adapting and extending the test suite based on observed behaviors. Importantly, achieving a 100% green test suite is often neither feasible nor the goal. Instead, the focus shifts to systematically reducing uncertainty and guiding the system toward acceptable behavior ranges.
Furthermore, during the design phase, it’s not uncommon that developers experiment with different base LLMs to evaluate their suitability for specific components. While traditional testing phases typically assume a fixed system configuration, in LLM-based development, variability in models, prompts, and hyperparameters is a key part of the design process. This means that testing is not simply a validation tool but a primary mechanism for system exploration and refinement.
This interplay between model selection, prompt engineering, and test design underscores the importance of integrating testing as an ongoing, foundational activity. The taxonomy we propose below supports this need by offering a structured framework for test case design that remains valuable across all stages of a system’s lifecycle–from early prototyping to deployment and long-term maintenance.
In this context, testing is not a final verification step but an integral driver of system development. This dynamic approach to TDD highlights that, for LLM-based systems, testing is even more critical than in traditional software development, where formal methods, type systems, and static analyses can partially guarantee behavior. In contrast, for LLM-centric systems, testing is often the primary means of ensuring reliability, safety, and functionality.
@@ 3. Taxonomy for LLM Test Case Design
3.1. Case: Issue report classification
We explain all four facets in support of a case SUT by name ClassifyIssueReport, which has the purpose to classify issue reports for issue tracking (such as on GitHub). Given an issue text, the SUT shall return exactly one label (for reasons of simplicity) - BUG, FEATURE, INVALID, or DUPLICATE. Invalid issue reports demand work without benefit (Laiq and Dobslaw, 2025). In case of a duplicate, i.e. an issue that has been brought up already in an existing issue report, point to the matching existing issue. The classification then could lead to different actions, for example assigning bugs to an appropriate tester or developer, or adding a duplication note in the thread with reference to a duplicate issue and closing it.
Classifying reported issues has practical relevance and can be solved in many different ways in support of traditional software, machine learning, or LLMs. Using LLM’s we could imagine simple solutions with a single model or agentic systems that collaborate over the entire projects issue reporting history to match potential duplicates.
@@ 3.2. Software Under Test
SUT Example Components. The ClassifyIssueReport SUT is a combination of a duplication checking function DuplicationFinder and an LLM call. For a given issue report, the function returns either a matching issue id - to then return DUPLICATE - or null. In the latter case the LLM is invoked. Model and Configuration. One LLM is used only when the duplication check returns null: Mistral-7B-v0.1. The DuplicateFinder tool is used in version 0.3. The LLM is called with a fixed prompt defining BUG, FEATURE, and INVALID and instructing to answer with exactly one of them. Result variability is limited using temperature=0.0, top_p=1.0, n=1, and max_tokens=16.
@@ 3.3. Goal
Goal Examples G1 Decision validity. Given one issue report, the system should return exactly one label from BUG, FEATURE, INVALID, DUPLICATE. P1.1 Single-label: the output contains exactly one label. P1.2 Membership: the label is strictly one of BUG, FEATURE, INVALID, DUPLICATE. P1.3 Format discipline: no extra content beyond the label. G2 Duplicate alignment. When the duplication function provides a matching id, the system should surface a DUPLICATE decision that references that id; when it provides no match, the system should decide among BUG, FEATURE, INVALID. P2.1 Tool consistency: if the duplication function returns an id, the decision is DUPLICATE and the same id is echoed. P2.2 Non-duplicate discipline: if the duplication function returns null, the decision is not DUPLICATE. P2.3 ID provenance: any reported duplicate id must come from the duplication function. G3 Consistent behavior. At fixed prompt/model/settings, the system should produce the same decision for the same input text. P3.1 Run-to-run stability: at fixed settings, the same input yields the same label. P3.2 Minor-text robustness: light, meaning-preserving paraphrases yield the same label. P3.3 Prompt-format robustness: benign formatting changes (e.g., whitespace, line breaks) do not change the label.
@@ 3.4. Oracles
Oracle Examples For G1 Decision validity. O1 (P1.1–P1.3) Per-case check: read the output; confirm it has exactly one label, that the label is in BUG, FEATURE, INVALID, DUPLICATE, and that nothing extra is included. O1-metric Dataset view: share of cases that pass O1. For G2 Duplicate alignment. O2 (P2.1) Per-case check when the tool returns an id: label is DUPLICATE and the same id is returned. O3 (P2.2) Per-case check when the tool returns null: label is not DUPLICATE. O2/O3-metric Dataset view: share of cases that follow the tool result (report separately for tool=id and tool=null). For G3 Consistent behavior. O4 (P3.1) Metric: repeatability — same label across re-runs with fixed settings. O5 (P3.2) Metric: paraphrase agreement — same label across light rewordings. O6 (P3.3) Metric: format agreement — same label across harmless formatting changes.
@@ 3.5. Inputs
[…]
4. Empirical Investigations
To demonstrate and test the practical value of our taxonomy, we conducted three complementary empirical investigations. Together, they examine how the taxonomy can guide both human and automated tool evaluation and how selected aspects of it manifest in practice. The first two studies focus on testing tools: Section 4.1 presents a manual mapping of six open-source frameworks, while Section 4.2 extends this analysis by turning the taxonomy into a detailed checklist that also serves as a structured prompt for LLM-based evaluation. This dual approach allowed us to assess how well the taxonomy captures existing practices and to explore whether LLMs can assist in identifying tool capabilities and gaps not yet represented in the taxonomy. The third study, in Section 4.3, shifts focus from tools to systems under test, applying a sensitivity analysis to a multi-agent Android testing system (Yoon et al., 2024) to illustrate how SUT variability appears across different LLM models and leads to output variability that motivates aggregated oracles. All supplementary materials, including tool documentation, analysis prompts, detailed results, and complete action logs, are available online [^1] https://doi.org/10.5281/zenodo.17393106.
4.1. Manual Taxonomy-Based Tool Evaluation
@@
We identified relevant GitHub repositories by searching for “LLM Testing” and “LLM Evaluation”, selecting frameworks that specifically target LLM applications (i.e., systems combining LLMs with prompt templates, optionally enhanced with retrieval mechanisms or tool-calling capabilities) and provide dedicated testing interfaces. The six selected projects, chosen based on GitHub popularity (stars) and distinct testing features, are Opik (14.4k), DeepEval (Team, 2024a) (11.4k stars), RAGAs (11k stars), Promptfoo (Team, 2024b) (8.6k stars), Phoenix (7.1k), and Giskard (Giskard Team, 2024) (4.9k stars) at the time of writing.
@@
Facet SUT Goal Oracles Inputs
Opik The tool supports visualization features for comparing different models and prompt versions. Testing goal can be set with one or more properties with varying number of inputs (dataset). Oracles are implemented using deterministic, context relevance, and LLM-as-a-judge metrics. Inputs with semantic variations can be synthesized (expanded from existing input examples) with an separate LLM agent and with custom variation instructions from users.
DeepEval The tool supports visualization features for comparing different models and prompt versions. Testing goal can be set with one or more properties for each test input. Oracles are implemented using deterministic and model-based metrics, including conversational/multi-modal support. Inputs with semantic variations can be synthesized with an separate LLM agent and with various resources (e.g., documents).
RAGAs The tool supports organizing responses from different models and prompt versions, while keeping the visualization features private Testing goal can be set with one or more properties with varying number of inputs (dataset). Oracles are implemented as metrics measuring context relevance, tool call accuracy, and similarity with reference answers either with llms or deterministic calculation. Inputs, specifically the queries for the retrieval-augmented systems, can be synthesized using a knowledge graph extracted from documents to ensure semantic diversity.
Promptfoo The tool supports visualization features for comparing different models and prompt versions. Testing goal can be set with one or more properties with varying number of inputs. Oracles are implemented using deterministic and model-based metrics, allowing weighted metric combinations. Inputs with semantic variations can be synthesized with an separate LLM agent.
Phoenix The tool supports visualization features for comparing different models and prompt versions. Testing goal can be set with one or more properties with varying number of inputs. Oracles are implemented using deterministic and LLM-as-a-judge metrics. Inputs can be synthesized with a custom prompt and an LLM.
Giskard The tool separates visualization and result comparison features into an external enterprise platform. Testing goal can be defined as one or more properties across a set of inputs. Oracles are typically implemented as dataset-wise scores via built-in or custom metric functions across multiple inputs. Inputs can be transformed using various predefined syntactic transformation methods.
@@
All tools accommodate variability across the three SUT sub-facets, though their capabilities differ slightly. Each allows users to specify the Component under test (e.g., a prompt template or agent), select the Model(s) (e.g., GPT-4o, Claude 3.7 Sonnet), and adjust Configurations (e.g., temperature, token limits) that influence runtime behavior. Notably, Promptfoo enables “unfolded” comparisons, allowing testers to assess differences between prompt versions, model choices, or configuration settings. Similarly, Opik and DeepEval include visualization tools but primarily focus on identifying regressions through aggregated metrics (e.g., the proportion of passing test cases within a dataset). While Giskard’s open-source SDK does not directly support cross-version SUT comparisons, this functionality is available through integration with an an external enterprise platform (Giskard, 2025).
@@
An ideal test case is defined with a clear and specific testing goal, typically composed of one or more relevant properties. In most frameworks, a test case consists of a single input paired with an expected output, evaluated against one or more property-checking metrics. Giskard places greater emphasis on dataset-level scores and uniquely offers fine-grained controls for selected properties in each test run, though such functionality is supported through the external enterprise platform. Certain LLM testing sub-goals—such as assessing robustness against prompt injection attacks—are best evaluated using a set of targeted inputs combined with specific properties of interest. While other tools can execute multiple test cases in a single evaluation, they lack explicit support for linking evaluations to a coherent test goal across different properties. Conversly, Giskard lacks the flexibility to configure individual test cases separately when needed.
@@
Evaluations typically yield either a continuous score (e.g., similarity to a ground-truth output on a 0-1 scale) or a binary pass/fail result, with continuous scores often thresholded into binary outcomes. Giskard mainly operates at the dataset level, aggregating individual assessments (e.g., reporting a 70% pass rate for politeness). However, LLM outputs can vary across repeated runs, and current tools lack aggregation methods to handle multiple evaluations of the same input within the same system under test (SUT). Instead, each run is either treated as an independent outcome, or simply averaged (Phoenix), ignoring stochastic variability.